### v2

# EvaLoop: Assessing LLM Robustness in Programming from a Self-consistency Perspective

Abstract
Assessing the programming capabilities of Large Language Models (LLMs) is crucial for their effective use in software engineering. Current evaluations, however, predominantly measure the accuracy of generated code on static benchmarks, neglecting the critical aspect of model robustness during programming tasks. While adversarial attacks offer insights on model robustness, their effectiveness is limited and evaluation could be constrained. Current adversarial attack methods for robustness evaluation yield inconsistent results, struggling to provide a unified evaluation across different LLMs. We introduce EvaLoop  [^1] The code is being organized, we will make it public as soon as possible., a novel assessment framework that evaluate the robustness from a self-consistency perspective, i.e., leveraging the natural duality inherent in popular software engineering tasks, e.g., code generation and code summarization. EvaLoop initiates a self-contained feedback loop: an LLM generates output (e.g., code) from an input (e.g., natural language specification), and then use the generated output as the input to produce a new output (e.g., summarizes that code into a new specification). EvaLoop repeats the process to assess the effectiveness of EvaLoop in each loop. This cyclical strategy intrinsically evaluates robustness without rely on any external attack setups, providing a unified metric to evaluate LLMs’ robustness in programming. We evaluate 16 prominent LLMs (e.g., GPT-4.1, O4-mini) on EvaLoop and found that EvaLoop typically induces a 5.01%-19.31% absolute drop in pass@1 performance within ten loops. Intriguingly, robustness does not always align with initial performance (i.e., one-time query); for instance, GPT-3.5-Turbo, despite superior initial code generation compared to DeepSeek-V2, demonstrated lower robustness over repeated evaluation loop. EvaLoop reveals distinct degradation patterns across different models, offering developers a simple-yet-effective assessment of how LLMs maintain functional coherence through multiple transformation loops. This is a critical factor for real-world software development, where requirements and implementations continuously evolve.
1 Introduction
Program synthesis using Large Language Models (LLMs) has emerged as a transformative approach in software engineering, with recent models demonstrating impressive capabilities in generating functionally correct programs from natural language descriptions chen2021evaluating ; roziere2023code ; nijkamp2023codegen ; lozhkov2024starcoder ; achiam2023gpt ; liu2024deepseek . As these models are increasingly integrated into development workflows mastropaolo2023robustness ; rasnayaka2024empirical ; jimenez2023swe , rigorously evaluating their performance becomes paramount to guarantee the reliability. However, current evaluations of LLM-based code generation predominantly focus on performance metrics such as pass@k, which measure success against predefined test suites chen2021evaluating ; austin2021program ; liu2024your . While valuable for gauging functional correctness on specific benchmark tasks, these metrics do not adequately address a critical dimension of generated code: robustness.
Initial prompt:
Write a python function to determine whether a given list of numbers forms a full sequence of
consecutive integers without any duplicates.
assert check_Consecutive([1,2,3,4,5]) == True
Solution:
def check_Consecutive(lst):
if not lst:
return False
s = set(lst)
return max(lst) - min(lst) + 1 == len(lst) and s == set(range(min(lst), max(lst) + 1))
[…]
Derived prompt:
Write a python function to determine if a list of numbers contains an unbroken sequence of consecutive
integers from its minimum to its maximum value.
assert check_Consecutive([1,2,3,4,5]) == True
Solution:
def check_Consecutive(lst):
if not lst:
return False
return set(lst) == set(range(min(lst), max(lst) + 1))
assert check_Consecutive([1,2,3,4,5]) == True
[…]
Figure 1: The generated solutions for MBPP+ No. 472 task by o3-mini. See full results in Appendix D.
Evaluating the robustness of LLMs in programming often involves adversarial attacks anand2021adversarial ; jha2023codeattack ; nguyen2023adversarial ; chen2023evaluating ; zhang2024attacks ; improta2025enhancing ; wei2023towards ; zaremba2025trading ; bavcic2024towards . In these methods, natural language prompts are perturbed at various levels (e.g., character, word, sentence), and the functional correctness of code generated from original versus perturbed prompts is compared to measure performance degradation, regarding as the robustness measurement. Although these adversarial attacks can quantify the robustness, their reliability and generalization face challenges. Our observations indicate that different adversarial attacks might "favor" different LLMs, potentially leading to contradictory conclusions about a model’s robustness depending on the specific attack employed. Specifically, when we applied three adversarial attacks—character, word, and sentence levels respectively—used in recode wei2023towards to the MBPP Plus benchmark liu2024your , and instructed GPT-3.5-Turbo to generate results for both original and adversarial prompts. As a result, GPT-3.5-Turbo only performs less robustly under word-level adversarial attack with synonym replacement. These findings suggest that a single adversarial attack may not provide a holistic view of an LLM’s robustness in programming, and a suite of diverse attacks might be necessary for a comprehensive assessment of a single model. However, this approach complicates the comparative evaluation of robustness across different models. For instance, if one model demonstrates strong robustness against attack X but performs poorly against attack Y, while another model shows the opposite pattern, it becomes challenging to definitively determine which model is inherently more robust overall. This motivates a question: Can we overcome the configuration-dependency of adversarial attacks to achieve a unified evaluation of LLMs’ robustness in programming?
Our solution. Inspired by the cognitive resilience associated with human mastery, we introduce EvaLoop, a novel framework for evaluating the LLMs’ robustness in programming. Echoing the Levels of Processing theory craik2002levels , deep cognitive engagement—such as that required to fully grasp dual relationships like language translation—builds robust knowledge structures inherently resistant to minor perturbations. EvaLoop operationalizes this concept by exploiting the duality wei2019code between complementary software engineering tasks (e.g., code generation li2022competition and code summarization ahmad2020transformer ) within a self-contained feedback loop. Given an initial natural language specification, an LLM generates code. This code is then summarized by the same LLM, creating a new specification for the subsequent evaluation loop. The feedback loop directly assesses the robustness of this self-consistency under iterative transformation, continuing until the generated code no longer passes functional tests. Hence, EvaLoop provides a uniform metric for LLMs’ robustness in programming—the average number of sustainable loops (ASL) before failure—eliminating dependence on external attack configurations and potential bias.
Figure 1 illustrates the performance of o3-mini on task No. 472 in MBPP Plus across several loops. Initially, as shown in Listing 1, o3-mini generates a correct solution that accurately checks if a list of numbers forms a complete sequence of consecutive integers without duplicates, handling both requirements. However, by the 4th loop (Listing 2), despite receiving a semantically equivalent prompt, o3-mini produces an incorrect solution. This later implementation fails to verify the absence of duplicates, only checking if the set formed from a range between the minimum and maximum values matches the set of the input list. Such an approach would erroneously return True for inputs with duplicates, demonstrating an inconsistency in the model’s reasoning across loops. This example demonstrates EvaLoop’s unique ability to deeply evaluate LLMs’ robustness in programming. By iteratively checking self-consistency through the dual tasks, it reveals subtle degradation in reasoning. Additionally, the capacity to deeply probe self-consistency across loops makes EvaLoop’s evaluation particularly relevant to practical software development, where understanding and measuring how code quality evolves over time is crucial.
Contributions. Our main contributions are:
  • Problem Identification: We identified the limitations of adversarial attack methods for evaluating LLMs’ robustness in programming that different adversarial attacks might “favor” different LLMs, potentially leading to biased measurement and exhibiting less reliability and generalization.
  • Novel Evaluation Framework (EvaLoop): We propose EvaLoop from a self-consistency perspective, which is based on a self-contained feedback loop and leverages duality between specific software engineering tasks. In this paper, we provide two type of self-contained feedback loop frameworks: (1) leveraging the inherent duality between code generation and code summarization, and (2) employing internal duality in code translation pan2023understanding ; yang2024exploring to build a code translation chain (e.g., Python → JavaScript → PHP → Ruby → Perl → Python). In both loop types, robustness is measured intrinsically by our proposed metric, i.e., ASL, which effectively eliminates the need for the dependency on external attack configurations and implicitly simulating aspects of software evolution under iterative stress.
  • Empirical Validation and Insights: Through empirical evaluation on 16 popular LLMs, we demonstrate that EvaLoop provides a stable and effective metric for quantifying programming robustness—an LLM’s ability to maintain functionally correct code through iterative modifications. Our findings indicate that applying EvaLoop typically leads to a 5.01%-19.31% absolute decrease in pass@1 performance within ten loops. Intriguingly, we observed that robustness does not necessarily correlate with initial code generation quality; for instance, GPT-3.5-Turbo chatgpt , despite superior initial performance compared to DeepSeek-Coder-V2-Lite-Instruct zhu2024deepseek , exhibited lower robustness (6.38 vs. 6.36) and final performance after evaluation by EvaLoop. Consequently, EvaLoop reveals distinct degradation patterns across LLMs, offering developers a preliminary assessment of how LLMs maintain functional coherence in different evaluation loops. This insight is practical for real-world software development, where requirements and implementations continuously evolve. Furthermore, EvaLoop also offers a novel perspective on the memorization versus comprehension debate in LLMs: the capacity to generate correct solutions over more loops may indicate a deeper comprehension to the prompt, rather than mere recall of training data.
2 Approach: EvaLoop
(a) EvaLoop: Code Generation-Summary loop.
(b) EvaLoop: Code Translation loop.
Figure 2: The overview of EvaLoop. ASL: Average number of sustainable loops.
Figure 2 presents an overview of EvaLoop. Drawing inspiration from the Levels of Processing theory craik2002levels , EvaLoop evaluates LLM robustness through iterative tasks requiring consistent semantic understanding—a self-consistency perspective. Our approach is highly flexible, straightforward to implement, and offers strong generalization capabilities. We leverage two software engineering dual relationships: the complementary nature of code generation and summarization, and the semantic preservation in code translation. As demonstrations, we implement two example frameworks: Figure 2(a) shows an evaluation loop integrating code generation and summarization, while Figure 2(b) presents a code translation chain across multiple programming languages, with each successful translation serving as a step in an extended evaluation loop.
2.1 Framework 1: Generation-Summarization Duality Loop
Figure 2(a) illustrates our first framework, which utilizes the intrinsic duality between code generation and code summarization tasks. This framework establishes a cyclical evaluation process designed to assess the LLMs’ robustness in programming in maintaining semantic consistency across multiple transformation loops.
The evaluation loop commences with an initial natural language prompt that specifies a programming task (e.g., "Write a Python function to…"). The process then unfolds as follows:
  1. 1.
    Code Generation Phase: The LLM under evaluation generates executable code based on the current prompt. The transformation from natural language to formal code evaluates the model’s ability to translate human intent into programmatically valid solutions.
  2. 2.
    Testing Phase: The generated code is executed rigiously against a predefined test suite to verify its functional correctness. This phase acts as a binary classifier, determining whether the solution satisfies the functionality requirements.
  3. 3.
    Bifurcation Point: Based on testing results, the process follows one of two paths: If the code passes all tests, the loop proceeds to the summarization phase. If the code fails any test, the loop terminates, and we record the number of loops the LLM reaches for this task.
  4. 4.
    Code Summarization Phase: Upon successful testing, the same LLM that generated the code is tasked with summarizing the implementation back into natural language. Such an inverse transformation tests the model’s comprehension of its own generated content.
  5. 5.
    Prompt Update and Evaluation: The model-generated summary serves as the updated prompt for the subsequent iteration, replacing the previous prompt. This creates a self-contained feedback loop where the model’s understanding of the task evolves with each loop.
The evaluation process continues iteratively until all code generated by the LLM fail the functional tests. The framework’s effectiveness is quantified using a robustness metric defined as the average number of evaluation loops the LLM can sustain before functional failure occurs across a comprehensive benchmark suite, i.e., ASL. A higher ASL indicates superior semantic consistency and robustness in the model’s understanding of both programming tasks and its own generated solutions.
The intuitive rationale behind this framework stems from a fundamental expectation: any intelligent system should comprehend what it produces. Just as humans generally understand their own statements and can explain their own work, a robust LLM should maintain semantic consistency when processing its own output. If a model generates code that it later cannot accurately summarize—or if it produces a summary that leads to functionally different code when used as a prompt—this reveals critical limitations in its semantic understanding. This framework provides a natural and intuitive test of a model’s coherence and stability by examining whether it can maintain the essential meaning of a programming task through successive transformations of its own creation. An LLM that quickly produces failing code after just a few loops demonstrates poor semantic stability, while models that sustain longer loops exhibit superior understanding of both the task and their own generated solutions.
2.2 Framework 2: Cyclical Code Translation Loop
Figure 2(b) illustrates the second framework, which evaluates LLMs’ robustness in programming through cyclical code translation across multiple programming languages. This framework rigorously tests the model’s ability to preserve semantic equivalence while traversing a chain of distinct programming paradigms and syntactic structures.
The evaluation process begins with functionally correct source code in an initial programming language (PL A in the Figure), which generally could be derived from a benchmark, e.g., MBPP Plus. The sequence unfolds as follows:
  1. 1.
    Initial Translation Prompt: The process commences with a prompt instructing the LLM to translate code from the source language (PL A) to a target language (PL B). For example, "Convert the following Java code to Python."
  2. 2.
    Code Translation Phase (PL A $\rightarrow$ PL B): The LLM performs the translation, generating equivalent code in the target language while preserving the original functionality.
  3. 3.
    Testing Phase: The translated code is executed thorough language-specific test suites designed to verify functional equivalence with the original implementation.
  4. 4.
    Bifurcation Point: Based on test results: If the code passes all tests, the process continues with an updated prompt for the next translation. If the code fails any test, the evaluation loop terminates, and we record the number of loops the LLM reaches for this task.
  5. 5.
    Translation Chain Progression: Upon successful testing, we update the previous prompt and instruct the LLM translate the code to the next language in the sequence (e.g., PL B $\rightarrow$ PL C). The process repeats, creating a chain of translations across multiple programming languages.
The evaluation loop can incorporate diverse programming languages with varying paradigms (e.g., imperative, functional, object-oriented) and eventually loop back to the original language. This creates a comprehensive test of the model’s cross-language understanding and its ability to maintain program semantics across multiple transformations. The framework’s effectiveness is also quantified using a robustness metric defined as the average number of successful translation steps the language model can sustain before producing functionally incorrect code, i.e., ASL (regard each step as a loop). A higher ASL superior semantic preservation capability across diverse programming ecosystems.
This framework rigorously assesse the LLMs’ robustness in programming when subjected to transformations across diverse programming languages. Such an capability is paramount, as code translation is not merely an academic exercise but a practical and increasingly frequent requirement for enterprises weisz2021perfection . Companies often need to migrate legacy systems or update their technology stack to meet new business demands or enhance system properties—for example, transitioning a project from Java to Rust to achieve greater memory safety and performance. Consequently, EvaLoop provides a critical methodology for gauging how well LLMs can support these complex, real-world software engineering endeavors, thereby reflecting their true translational and adaptive capabilities.
3 Evaluation
Experimental Setup. Our evaluation employs pass@k roziere2023code to assess the functional correctness of LLM-generated code at each evaluation loop. For practical relevance, we set k to 1, i.e., pass@1, reflecting real-world deployment scenarios where code assistants such as GitHub Copilot only returns one solution to a one-time query from end-users. Our evaluation examines two distinct inference strategies: greedy decoding (temperature = 0) and temperature sampling renze2024effect (temperature = 0.2), following established best practices from prior research liu2024your to optimize the performance-diversity tradeoff in code generation. For comprehensive robustness assessment, we compute a new metric, the average number of sustainable loops (ASL) until functional failure observed across all benchmark coding tasks. It could be computed as follows:
$\text{ASL}=\frac{\sum_{i=1}^{M}i\times n_{i}}{N},$ (1)
where $i$ means the loop index [^2] In code translation, per programming language translation is a loop, $n_{i}$ represents the number of tasks that passed only at loop $i$, and $N$ represents the total number of tasks. To ensure generalizability and representativeness, we conducted an extensive evaluation across 16 state-of-the-art LLMs, encompassing both proprietary closed-source models and open-source alternatives. See details in Appendix A.
Benchmark Selection. EvaLoop is designed for versatility and seamless integration with existing programming benchmarks. In this study, we utilize two comprehensive benchmark suites, MBPP Plus liu2024your and MXEVAL athiwaratkun2022multi , which offer significant advancements over their predecessors. MBPP Plus enhances the original MBPP austin2021program by applying rigorous quality filtering to programming tasks and expanding the test case corpus 35-fold. MXEVAL extends MBPP’s Python-centric focus to encompass 12 distinct programming languages. For Framework 1 (Generation-Summarization Duality Loop), we evaluate the LLMs’ robustness in programming on 378 distinct tasks from MBPP Plus. The operational loop for this framework is limited to a maximum of 10 loops. For Framework 2 (Cyclical Code Translation Loop), we establish a circular code translation chain: Python → PHP → Ruby → JavaScript → Perl → Python. The four intermediate languages were randomly selected from the non-Python languages available in MXEVAL [^3] Our experimental results showed no significant differences between two such distinct random selections.. This cyclical translation process utilizes a carefully curated subset of 375 tasks from the multilingual MXEVAL.
Figure 3: Distribution of average sustainable cycles per task in MBPP Plus. The data integrates results from greedy decoding and temperature sampling across 16 popular LLMs (sorted from smallest to largest according to the results of greedy decoding).
Figure 4: Robustness assessment of LLMs on Framework 1. See their full names in Appendix A.
Greedy Decoding vs. Temperature. Figure 3 delineates the performance distribution, measured in ASL per task on the MBPP Plus benchmark, across 16 leading LLMs employing either greedy decoding or temperature-controlled stochastic sampling. The analysis reveals a counter-intuitive finding: greedy decoding achieves a marginally superior average performance (6.64 loops) over temperature sampling (6.48 loops). This outcome is particularly noteworthy given that temperature sampling is often the pre-selected decoding mechanism in numerous LLMs (e.g., ChatGPT employs a default temperature of 0.7), primarily valued for its capacity to introduce variability and emulate human-like expression. A plausible explanation for this finding is that the advantages of temperature sampling, such as improved "naturalness" in textual output, are less critical in the domain of code generation. Code inherently demands strict adherence to precise semantic and syntactic rules, a context in which the deterministic, highest-probability path selection of greedy decoding may offer a more consistent advantage over the stochastic nature of temperature sampling. This finding motivates the need for further research into optimal decoding strategy selection specifically tailored for the domain of code generation. Based on the results, we only focus on the results of greedy encoding in the rest of the experiments.
Effectiveness of Framework 1. Figure 4 illustrates the pass@1 rate with greedy decoding across evaluation loops on MBPP Plus. Our analysis reveals that after multiple evaluation loops, all LLMs’ performance degradation by a large margin, absolutely ranging from 5.01% to 19.31% compared to their initial performance. Prominent open-source LLMs such as LLaMA-3.1-8B (lm8b) grattafiori2024llama , Codestral-22B (cs22b) mistral2024codestral , and OpenCoder-8B (oc8b) huang2024opencoder exhibit absolute performance reductions of 19.31%, 6.35%, and 10.58%, respectively. This pattern extends to closed-source models as well, with ChatGPT family members—including o4-mini (one of the most powerful LLMs available), o1-mini, and GPT-4-Turbo—showing absolute performance decreases of 10.05%, 10.58%, and 6.61%, respectively. These findings substantiate that EvaLoop can reveal distinct degradation patterns across different models, providing developers with a more authentic assessment of how LLMs maintain functional coherence through multiple iterations. Consequently, EvaLoop serves as a valuable complement to traditional evaluation methods, offering deeper insights into the LLMs’ robustness in programming in realistic software development environments.
Beyond the performance degration, EvaLoop also uncovers notable re-rankings of LLM performance relative to initial accuracy (i.e., results at loop 1). Figure 4 demonstrates that while OpenCoder-8B (oc8b) initially outperforms Codestral-22B (cs22b), this advantage reverses after the first evaluation loop, with Codestral-22B ultimately demonstrating superior sustained performance. A similar pattern emerges among closed-source LLMs: despite o4-mini achieving the highest initial performance, models such as o3-mini, gpt-4.1-mini, and Qwen2.5-Coder-32B hui2024qwen2 exhibit better performance in later loops. Particularly noteworthy is Qwen2.5-Coder-32B, an open-source model that initially underperforms several ChatGPT family members but ultimately demonstrates the highest programming robustness. This suggests that open-source LLMs with tailored design and training methodologies can match or even surpass closed-source alternatives in maintaining performance integrity through iterative transformations. In summary, our framework provides a fine-grained, comprehensive, and ecologically valid evaluation methodology for assessing LLMs’ robustness in programming in real-world scenarios.
Examining ASL metric reveals that o3-mini achieves the best performance not only on the initial accuracy (80.95% at pass@1) but also consistent robustness (ASL with 7.58) throughout each evaluation loop. While Qwen2.5-Coder-32B exhibits superior performance in the last three evaluation loops, its comparatively modest initial accuracy results in a slightly lower ASL value with 7.42. This underscores how our proposed ASL metric effectively balances performance across all evaluation stages, providing a more comprehensive assessment than single-point measurements. Particularly noteworthy is LLaMA-3.1-8B, which consistently underperforms across all metrics at all evaluation points—especially in ASL, where it scores (below 5) substantially lower than other LLMs. This pronounced disparity highlights the significant challenges some models face in maintaining robustness throughout iterative evaluation loops, reinforcing the value of our framework in identifying models with sustained performance integrity versus those that degrade substantially under realistic usage conditions.
Figure 5: Robustness assessment of LLMs on Framework 2.
Effectiveness of Framework 2. Figure 5 illustrates the pass@1 rate with greedy decoding across successive evaluation loops that involve a sequential chain of code translations. The baseline in x-axis presented in the figure indicates the initial accuracy of each LLM on the MBPP Plus (Python) benchmark. Our analysis reveals a noteworthy trend: consistent with observations from Framework 1, all LLMs exhibit a substantial degradation in performance upon completion of the entire code translation cascade. This performance decline, absolutely ranging from 32.00% to 54.13%, compared to their initial baseline performance. Specifically, while CodeQwen1.5-7B hui2024qwen2 initially outperforms OpenCoder-8B, deepseek-coder-7b guo2024deepseek , and Codestral-22B, the latter models demonstrate greater robustness following the chain of programming language translations. A similar pattern emerges among closed-source LLMs. For instance, GPT-4.1, despite an initial performance deficit compared to o4-mini, o3-mini, and GPT-4.1-mini, ultimately achieves the highest performance after evaluation by EvaLoop. Generally, this re-ranking phenomenon mirrors the outcomes observed in Framework 1.
Examining ASL metric reveals another insight into LLMs’ robustness in programming. GPT-4.1 leads with an ASL of 3.42, closely followed by GPT-4.1-mini (3.40) and o3-mini (3.40), signifying not only strong initial accuracy but also remarkable performance consistency across sequential translation tasks, underscoring their robust nature and the superior. In stark contrast, open-source LLMs generally represent a relatively low ASL value, and LLaMA-3.1-8B records the lowest ASL with 2.11, whose trajectory shows a low baseline followed by continuous, pronounced degradation, highlighting difficulties in maintaining functional integrity. ASL is invaluable for a holistic assessment, moving beyond single-point evaluations to capture sustained operational efficacy and distinguish models with enduring performance integrity from those that falter under the cumulative stress of multi-step, cross-lingual code translations.
Memorization vs. Comprehension Our proposed metric, ASL, potentially offers another lens through which we can explore the nature of an LLM’s success—differentiating between apparent memorization schwarzschild2024rethinking ; wang2024generalization and deeper conceptual comprehension. High value of ASL (visible as higher bars towards the right of Figure 3) indicates that the LLM carries a more robust understanding of the underlying problem, encompassing both the provided code and its corresponding natural language prompt. In our evaluation framework, each loop potentially involves variations in the generated code and its subsequent summarization. To consistently maintain correctness across these extended evaluation sequences, the model must possess a profound grasp of the specific task’s semantics and logic, rather than merely matching superficial patterns. This capacity to adapt and retain accuracy despite iterative semantic-equivalent changes.
Conversely, tasks exhibiting low ASL, particularly those around 1.x or 0.x (represented by lower bars towards the left of Figure 3), paint a different picture. Such scores signify that the model, while potentially succeeding in the initial evaluation loop (thereby achieving an initial accuracy), fails to maintain correctness in subsequent iterations. This rapid degradation in performance could imply that the initial success was not rooted in genuine understanding. Instead, it might suggest that the model recognized and reproduced patterns or specific examples encountered during its training phase introduced by potential data leakage deng2023benchmark ; balloccu2024leak , building only a shallow memorization. When faced with the slightly altered context of a new summarization or the need to iterate upon its own generated code in the next loop, the model’s lack of foundational understanding leads to the generation of an incorrect solution. Thus, the iterative evaluation of sustainable loops serves as a valuable diagnostic for probing whether an LLM is truly comprehending tasks or relying on learned exemplars or potential data leakage.
4 Related Work
LLMs’ Robustness in Programming. Code generation has emerged as the predominant practical application of AI in programming environments liang2024large , catalyzing extensive research into the robustness of LLMs in this domain. Empirical assessments of commercial systems reveal concerning vulnerabilities: GitHub Copilot produces inconsistent implementations for semantically equivalent prompts in 46% of cases, with correctness variations reaching 28% between prompt variants mastropaolo2023robustness . Similarly, Codex exhibits substantial sensitivity to minor linguistic perturbations in task specifications shirafuji2023exploring ; zhuo2023robustness . The research community has responded with LLMs’ robustness through diverse adversarial techniques chen2023evaluating ; zhang2024attacks ; jha2023codeattack ; bavcic2024jy61 alongside corresponding defensive strategies improta2025enhancing ; yang2024robustness ; ge2024demonstration . However, different adversarial attacks might "favor" different LLMs, potentially leading to contradictory conclusions about a model’s robustness depending on the specific attack employed, struggling in a unified evaluation. EvaLoop addresses this critical gap by leveraging the duality between generation and understanding, enabling unified robustness assessment across LLMs with the average number of sustainable loops, i.e., ASL . This novel approach provides a natural measure of LLMs’ robustness in programming.
Self-consistency in LLMs. Self-consistency in LLMs, originally introduced by Wang et al., represents a novel decoding strategy that significantly enhances reasoning accuracy by sampling diverse reasoning paths and selecting the most consistent answer through majority voting. Beyond using self-consistency to improve reasoning capability, researchers have begun exploring it as an evaluation metric for LLM performance min2023beyond ; lee2024evaluating .
IdentityChain min2023beyond is the most closely related work, employing a feedback loop between natural language specifications and code generation to assess semantic understanding in Code LLMs. However, IdentityChain fundamentally differs from EvaLoop in both objectives and methodology. While IdentityChain evaluates whether models can maintain semantic coherence in a single round-trip translation between natural language and code, EvaLoop addresses a distinct challenge: evaluating robustness through sustained iterative transformations that mirror real-world software evolution.
EvaLoop introduces several key innovations that distinguish it from IdentityChain: (1) a cross-language translation chain that stress-tests models across multiple programming paradigms rather than a single NL-to-code-to-NL cycle, (2) the Average Sustainable Loops (ASL) metric that quantifies long-term stability under continuous transformation pressure rather than one-time accuracy, and (3) a focus on functional correctness preservation through multiple iterations rather than semantic equivalence in a single translation. This fundamental difference in scope—from IdentityChain’s semantic consistency to EvaLoop’s robustness under iterative stress—makes EvaLoop uniquely suited for evaluating how Code LLMs perform in dynamic, evolving software development scenarios where code undergoes repeated modifications and translations.
5 Conclusion
In this paper, we addressed the critical gap in current evaluations of LLMs’ programming capabilities, which often overlook robustness in favor of initial accuracy. To this, we introduced EvaLoop, a novel framework that assesses LLMs’ robustness in programming from a self-consistency perspective, i.e., leveraging the duality of software engineering tasks, such as code generation and summarization. EvaLoop employs a self-contained feedback loop, iteratively transforming code or its natural language specification until functional degradation occurs, thereby achieving unified evaluation to LLMs’ robustness without relying on external adversarial configurations. Our empirical evaluation across 16 prominent LLMs demonstrated that EvaLoop with framework 1 effectively reveals distinct degradation patterns, typically inducing a 5.01%-19.31% absolute drop in pass@1 performance within ten loops, and importantly, showed that initial programming proficiency does not always correlate with sustained robustness. Hence, EvaLoop provides a more authentic and nuanced analysis of how LLMs maintain functional coherence under iterative stress, offering unique insights for their practical deployment in dynamic software development environments and contributing to the broader discussion on model understanding versus memorization.
Looking ahead, we plan to extend the application of EvaLoop to a broader range of LLMs and diverse programming benchmarks. This expansion aims to further validate our approach and contribute towards establishing a more unified and comprehensive standard for evaluating the iterative robustness of code-generating models.
Appendix A LLM Names Mapping
Table 1: Mapping of short model names to their full names or known identifiers.
Short Name Full Name / Identifier
cs22b Codestral-22B-v0.1 [30]
ds7b deepseek-coder-7b-instruct-v1.5 [16]
dsv2 DeepSeek-Coder-V2-Lite-Instruct [55]
lm8b Llama-3.1-8B-Instruct [15]
oc8b OpenCoder-8B-Instruct [17]
qw32b Qwen2.5-Coder-32B-Instruct [18]
qw7b CodeQwen1.5-7B-Chat [18]
gpt-3.5-turbo gpt-3.5-turbo-0125 [33]
gpt-4.1 gpt-4.1-2025-04-14 [38]
gpt-4.1-mini gpt-4.1-mini-2025-04-14 [39]
gpt-4o gpt-4o-2024-08-06 [35]
gpt-4-turbo gpt-4-turbo-2024-04-09 [34]
o1 o1-2024-12-17 [36]
o1-mini o1-mini-2024-09-12 [37]
o3-mini o3-mini-2025-01-31 [40]
o4-mini o4-mini-2025-04-16 [41]
Appendix B Limitation
EvaLoop provides significant insights into LLM programming robustness, yet several limitations warrant discussion. Primarily, our assessment relies exclusively on MBPP-derived benchmarks (MBPP Plus and MXEVAL). Despite their enhancements—rigorous quality filtering and expanded test cases—these benchmarks may not fully mirror the diversity of real-world programming tasks. We plan to extend EvaLoop to other benchmark families like HumanEval, APPS, and domain-specific datasets to validate the broader applicability of our findings. Furthermore, computational limitations constrained our evaluation to 16 LLMs, preventing a comprehensive survey of all existing models. Although we carefully selected a mix of leading open and closed-source models based on their MBPP Plus performance and resource accessibility, incorporating additional models, especially novel architectures or specialized coding LLMs, could reveal more nuanced robustness patterns throughout the LLM ecosystem.
Appendix C Experimental Environment
We conducted all experiments on a Ubuntu server with an AMD Ryzen Threadripper PRO 7955WX CPU, 256 GB RAM, and dual NVIDIA A6000 GPUs.
Appendix D Full Loop-Trace Examples for Frameworks 1 & 2
This section presents four complete loop traces. Figure 6 and Figures 78 illustrate two examples for Framework 1. For Framework 2, Figure 9 illustrates the case that the LLM successfully passes the evaluation and generates correct translations for each programming language. Meanwhile, Figure 10 highlights a borderline failure case: the LLM could generate initially correct Python solution while generating wrong python solution after a transaltion chain.
Figure 6: Full code and description generated by o3-mini for task 472 in MBPP Plus.
Figure 7: Full code and description generated by o4-mini for task No. 781 in MBPP Plus (Loop 1-5).
Figure 8: Full code and description generated by o4-mini for task No. 781 in MBPP Plus (Loop 6-9).
Figure 9: Full code translation results generated by gpt-4.1 for task No. 19 in MBPP Plus.
Figure 10: Full code translation results generated by gpt-4.1 for task No. 604 in MBPP Plus.
Appendix E Extra Result Exhibition for Framework 2
In Figure 11, we present that how many tasks pass (pass@1) for a programming language without considering the test results of the previous languages. Although the final pass rate has relatively improved, but these is also an obvious performance degradation, keeping consistent with Figure 5.
Figure 11: Original pass counts per language in Framework 2.